OpenStudy (gagecool):

Evaluate the function for x = 2: f(x) = 7x - 5 is A. 9 B. 19 C. -9 D. -19 Need help ASAP. Will give medal

OpenStudy (gagecool):

will give out medal to the person who give me the answer and explains to me why its that answer

OpenStudy (qwertty123):

f(x)= 7x-5 so f(2)=7*2-5=?

OpenStudy (qwertty123):

Hope this helps you @Gagecool

OpenStudy (xmissalycatx):

What you're going to want to do is plug in the "x-value" that it gave you into your equation.. So you'll plug 2 into the equation which gives you f(x) = 7 (2) -5 From there you use PEMDAS to solve the rest.. So multiply 7 and 2 which is 14, then subtract 5 from it when gives you __?__

OpenStudy (qwertty123):

7 times 2-5=?

OpenStudy (qwertty123):

@Gagecool What did you get?